# Chain-linked Multiple Matrix Integration via Embedding Alignment

### v2

@@ 4.1 Estimation error of CMMI
We simulate a chain of $(L+1)$ overlapping observed submatrices $\{\mathbf{A}^{(i)}\}_{i=0}^{L}$ for the underlying population matrix $\mathbf{P}$ as described in Figure 3, and then predict the yellow unknown block by Algorithm 1. Each $\mathbf{P}^{(i)}$ has the same dimension, i.e. $n_{i}\equiv n=pN$ for all $i=0,1,\dots,L$, and the overlap between $\mathbf{P}^{(i-1)}$ and $\mathbf{P}^{(i)}$ are set to $n_{i-1,i}\equiv m=\breve{p}n$ for all $i=1,\dots,L$. We generate $\mathbf{P}=\mathbf{U}\bm{\Lambda}\mathbf{U}^{\top}$ by sampling $\mathbf{U}$ uniformly at random from the set of $N\times 3$ matrices with orthonormal columns and set $\bm{\Lambda}=\text{diag}(N,\tfrac{3}{4}N,\tfrac{1}{2}N)$, so that $\lambda_{\min}=\tfrac{1}{2}N$ in this setting. We then generate symmetric noise matrices $\{\mathbf{N}^{(i)}\}$ with $\mathbf{N}^{(i)}_{st}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\sigma^{2})$ for all $i=0,1,\dots,L$ and all $s,t\in[n]$ with $s\leq t$. Finally we set $\mathbf{A}^{(i)}=(\mathbf{P}^{(i)}+\mathbf{N}^{(i)})\circ\bm{\Omega}^{(i)}$ where $\bm{\Omega}^{(i)}$ is a symmetric $n\times n$ matrix whose upper triangular entries are i.i.d. Bernoulli random variables with success probability $q_{i}\equiv q$.
@@
Recall that, by Theorem 2, the estimation error for $\hat{\mathbf{P}}_{\mathcal{U}_{0},\mathcal{U}_{L}}-\mathbf{P}_{\mathcal{U}_{0},\mathcal{U}_{L}}$can be decomposed into the first order approximation $\mathbf{M}_{\star}:=\mathbf{E}^{(1)}\mathbf{X}_{\mathcal{U}_{1}}(\mathbf{X}_{\mathcal{U}_{1}}^{\top}\mathbf{X}_{\mathcal{U}_{1}})^{-1}\mathbf{X}_{\mathcal{U}_{L}}^{\top}+\mathbf{X}_{\mathcal{U}_{1}}(\mathbf{X}_{\mathcal{U}_{L}}^{\top}\mathbf{X}_{\mathcal{U}_{L}})^{-1}\mathbf{X}_{\mathcal{U}_{L}}^{\top}\mathbf{E}^{({L})}$ and the remainder term $\mathbf{R}^{(0,L)}+\mathbf{S}^{(0,1,\dots,L)}$. Furthermore we also have
@@
$\|\mathbf{M}_{\star}\|_{\max}\lesssim\frac{(1+\sigma)\log^{1/2}n}{(pqN)^{1/2}},\quad\|\mathbf{R}^{(0,L)}+\mathbf{S}^{(0,1,\dots,L)}\|_{\max}\lesssim L\Big{(}\frac{(1+\sigma)^{2}\log n}{pqN}+\frac{(1+\sigma)\log^{1/2}n}{q^{1/2}\breve{p}^{1/2}pN}\Big{)}$ (4.1)
@@
with high probability. We compare the error rates for $\|\hat{\mathbf{P}}_{\mathcal{U}_{0},\mathcal{U}_{L}}-\mathbf{P}_{\mathcal{U}_{0},\mathcal{U}_{L}}\|_{\max}$ against $\|\mathbf{M}_{\star}\|_{\max}$ and $\|\mathbf{R}^{(0,L)}+\mathbf{S}^{(0,1,\dots,L)}\|_{\max}$ by varying the value of one parameter among $N$, $p$, $\breve{p}$, $q$, $L$ and $\sigma$ while fixing the values of the remaining parameters. Empirical results for these error rates, averaged over $100$ Monte Carlo replicates, are summarized in Figure 4. We note that the error rates in Figure 4 are consistent with the bounds in Eq. (4.1) obtained by Theorem 2.
@@
Figure 4: Log-log plot of the empirical error rates for $\|\hat{\mathbf{P}}_{\mathcal{U}_{0},\mathcal{U}_{L}}-\mathbf{P}_{\mathcal{U}_{0},\mathcal{U}_{L}}\|_{\max}$ (blue lines), its first order approximation $\|\mathbf{M}_{\star}\|_{\max}$ (orange lines), and the remainder term $\|\mathbf{R}^{(0,L)}+\mathbf{S}^{(0,1,\dots,L)}\|_{\max}$ (green lines) as we vary the value of a single parameter among $\{N,p,\breve{p},q,L,\sigma\}$ while keeping the values of the remaining parameters fixed (note that we set $\lambda_{\min}=\tfrac{1}{2}N$). Panel (a): vary $N\in\{300,600,1200,2400,4800,6000\}$ for $p=0.3,\breve{p}=0.1,q=0.8,L=2,\sigma=0.5$. Panel (b): vary $p\in\{0.05,0.1,0.15,0.2,0.25,0.35\}$ for $N=2400$, $\breve{p}=0.1$, $q=0.8$, $L=2$, $\sigma=0.5$. Panel (c): vary $\breve{p}\in\{0.1,0.2,0.4,0.6,0.8\}$ for $N=2400$, $p=0.3$, $q=0.8$, $L=2$, $\sigma=0.5$. Panel (d): vary $q\in\{0.1,0.2,\dots,0.9,1\}$ for $N=2400$, $p=0.3$, $\breve{p}=0.1$, $L=2$, $\sigma=0.5$. Panel (e): vary $L\in\{1,2,3,4,5,6\}$ for $N=2400$, $p=0.15$, $\breve{p}=0.1$, $q=0.8$, $\sigma=0.5$. Panel (f): vary $\sigma\in\{0.3,0.5,1,3,5,10\}$ for $N=2400$, $p=0.15$, $\breve{p}=0.1$, $q=0.8$, $L=2$, $\sigma=0.5$. Error rates in each panel are averages based on $100$ independent Monte Carlo replicates.
We next compare the entrywise behavior of $\hat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}-\mathbf{P}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}$ against the limiting distributions in Theorem 3. In particular, we plot in Figure 5 histograms (based on $1000$ independent Monte Carlo replicates) of the $(i,j)$th entries where $(i,j)\in\{(1,1),(1,2),(1,3)\}$, and it is clear that the empirical distributions in Figure 5 are well approximated by the normal distributions with parameters given in Theorem 3.
Figure 5: Histograms of the empirical distributions for the $(i,j)$th entry of $\hat{\mathbf{P}}_{\mathcal{U}_{0},\mathcal{U}_{L}}-\mathbf{P}_{\mathcal{U}_{0},\mathcal{U}_{L}}$; here $(i,j)=(1,1)$ (left panel), $(i,j)=(1,2)$ (middle panel), and $(i,j)=(1,3)$ (right panel). These histograms are based on $1000$ independent Monte Carlo replicates where $N=6000$, $p=0.3$, $\breve{p}=0.1$, $q=0.8$, $L=2$, and $\sigma=0.5$. The red lines are pdfs of the normal distributions with parameters given in Theorem 3.
@@ 4.2 Comparison with other matrix completion algorithms
We vary $L$ and compare the performance of Algorithm 1 (CMMI) with some existing state-of-art low-rank matrix completion algorithms, including generalized spectral regularization (GSR) (Mazumder et al., 2010), fast alternating least squares (FALS) (Hastie et al., 2015), singular value thresholding (SVT) (Cai et al., 2010), universal singular value thresholding (USVT) (Chatterjee, 2015), iterative regression against right singular vectors (IRRSV) (Troyanskaya et al., 2001). Note that increasing $L$ leads to more observed submatrices but, as each submatrix is of smaller dimensions, the total number of observed entries decreases with $L$ at rate of $N^{2}q/L$. Our performance metric for recovering the yellow unknown block is in terms of the relative Frobenius norm error $\|\hat{\mathbf{P}}_{\mathcal{U}_{0},\mathcal{U}_{L}}-\mathbf{P}_{\mathcal{U}_{0},\mathcal{U}_{L}}\|_{F}/\|\mathbf{P}_{\mathcal{U}_{0},\mathcal{U}_{L}}\|_{F}$. Plots of the error rates (averaged over $100$ independent Monte Carlo replicates) for different algorithms and their running times are presented in the left and right panels of Figure 6, respectively. Figure 6 shows that CMMI outperforms all competing methods in terms of both recovery accuracy and computational efficiency.

### v4

@@ 4.1 Estimation error of CMMI
We simulate a chain of $(L+1)$ overlapping observed submatrices $\{\mathbf{A}^{(i)}\}_{i=0}^{L}$ for the underlying population matrix $\mathbf{P}$ as described in Figure 3, and then predict the unknown yellow block by Algorithm 1. Each $\mathbf{P}^{(i)}$ has the same dimension, i.e. $n_{i}\equiv n=pN$ for all $i=0,1,\dots,L$, and the overlap between $\mathbf{P}^{(i-1)}$ and $\mathbf{P}^{(i)}$ are set to $n_{i-1,i}\equiv m=\breve{p}n$ for all $i=1,\dots,L$. We generate $\mathbf{P}=\mathbf{U}\bm{\Lambda}\mathbf{U}^{\top}$ by sampling $\mathbf{U}$ uniformly at random from the set of $N\times 3$ matrices with orthonormal columns and set $\bm{\Lambda}=\text{diag}(N,\tfrac{3}{4}N,\tfrac{1}{2}N)$. We then generate symmetric noise matrices $\{\mathbf{N}^{(i)}\}$ with $\mathbf{N}^{(i)}_{st}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\sigma^{2})$ for all $i=0,1,\dots,L$ and all $s,t\in[n]$ with $s\leq t$. Finally, we set $\mathbf{A}^{(i)}=(\mathbf{P}^{(i)}+\mathbf{N}^{(i)})\circ\bm{\Omega}^{(i)}$ where $\bm{\Omega}^{(i)}$ is a symmetric $n\times n$ matrix whose upper triangular entries are i.i.d. Bernoulli random variables with success probability $q_{i}\equiv q$.
@@
Recall that, by Theorem 2, the estimation error for $\widehat{\mathbf{P}}_{\mathcal{U}_{0},\mathcal{U}_{L}}-\mathbf{P}_{\mathcal{U}_{0},\mathcal{U}_{L}}$can be decomposed into the first order approximation $\mathbf{M}_{\star}:=\mathbf{E}^{(1)}\mathbf{X}_{\mathcal{U}_{1}}(\mathbf{X}_{\mathcal{U}_{1}}^{\top}\mathbf{X}_{\mathcal{U}_{1}})^{-1}\mathbf{X}_{\mathcal{U}_{L}}^{\top}+\mathbf{X}_{\mathcal{U}_{1}}(\mathbf{X}_{\mathcal{U}_{L}}^{\top}\mathbf{X}_{\mathcal{U}_{L}})^{-1}\mathbf{X}_{\mathcal{U}_{L}}^{\top}\mathbf{E}^{({L})}$ and the remainder term $\mathbf{R}^{(0,L)}+\mathbf{S}^{(0,1,\dots,L)}$. Furthermore we also have
@@
$\|\mathbf{M}_{\star}\|_{\max}\lesssim\frac{(1+\sigma)\log^{1/2}n}{(pqN)^{1/2}},\quad\|\mathbf{R}^{(0,L)}+\mathbf{S}^{(0,1,\dots,L)}\|_{\max}\lesssim L\Big{(}\frac{(1+\sigma)^{2}\log n}{pqN}+\frac{(1+\sigma)\log^{1/2}n}{q^{1/2}\breve{p}^{1/2}pN}\Big{)}$
@@
with high probability. We compare the error rates for $\|\widehat{\mathbf{P}}_{\mathcal{U}_{0},\mathcal{U}_{L}}-\mathbf{P}_{\mathcal{U}_{0},\mathcal{U}_{L}}\|_{\max}$ against $\|\mathbf{M}_{\star}\|_{\max}$ and $\|\mathbf{R}^{(0,L)}+\mathbf{S}^{(0,1,\dots,L)}\|_{\max}$ by varying the value of one parameter among $N$, $p$, $\breve{p}$, $q$, $L$ and $\sigma$ while fixing the values of the remaining parameters. Empirical results for these error rates, averaged over $100$ Monte Carlo replicates, are summarized in Figure 4. We note that the error rates in Figure 4 are consistent with the above bounds.
@@
Figure 4: Log-log plot of the empirical error rates for $\|\widehat{\mathbf{P}}_{\mathcal{U}_{0},\mathcal{U}_{L}}-\mathbf{P}_{\mathcal{U}_{0},\mathcal{U}_{L}}\|_{\max}$ (blue lines), its first order approximation $\|\mathbf{M}_{\star}\|_{\max}$ (orange lines), and the remainder term $\|\mathbf{R}^{(0,L)}+\mathbf{S}^{(0,1,\dots,L)}\|_{\max}$ (green lines) as we vary the value of a single parameter among $\{N,p,\breve{p},q,L,\sigma\}$ while keeping the values of the remaining parameters fixed (note that we set $\lambda_{\min}=\tfrac{1}{2}N$). Panel (a): vary $N\in\{300,600,1200,2400,4800,6000\}$ for $p=0.3,\breve{p}=0.1,q=0.8,L=2,\sigma=0.5$. Panel (b): vary $p\in\{0.05,0.1,0.15,0.2,0.25,0.35\}$ for $N=2400$, $\breve{p}=0.1$, $q=0.8$, $L=2$, $\sigma=0.5$. Panel (c): vary $\breve{p}\in\{0.1,0.2,0.4,0.6,0.8\}$ for $N=2400$, $p=0.3$, $q=0.8$, $L=2$, $\sigma=0.5$. Panel (d): vary $q\in\{0.1,0.2,\dots,0.9,1\}$ for $N=2400$, $p=0.3$, $\breve{p}=0.1$, $L=2$, $\sigma=0.5$. Panel (e): vary $L\in\{1,2,3,4,5,6\}$ for $N=2400$, $p=0.15$, $\breve{p}=0.1$, $q=0.8$, $\sigma=0.5$. Panel (f): vary $\sigma\in\{0.3,0.5,1,3,5,10\}$ for $N=2400$, $p=0.15$, $\breve{p}=0.1$, $q=0.8$, $L=2$, $\sigma=0.5$. Error rates in each panel are averages based on $100$ independent Monte Carlo replicates.
@@ 4.2 Comparison with other matrix completion algorithms
We vary $L$ and compare the performance of Algorithm 1 (CMMI) with some existing state-of-art low-rank matrix completion algorithms, including generalized spectral regularization (GSR) (Mazumder et al., 2010), fast alternating least squares (FALS) (Hastie et al., 2015), singular value thresholding (SVT) (Cai et al., 2010), universal singular value thresholding (USVT) (Chatterjee, 2015), iterative regression against right singular vectors (IRRSV) (Troyanskaya et al., 2001). Note that increasing $L$ leads to more observed submatrices but, as each submatrix is of smaller dimensions, the total number of observed entries decreases with $L$ at rate of $N^{2}q/L$. Our performance metric for recovering the yellow unknown block is in terms of the relative Frobenius norm error $\|\widehat{\mathbf{P}}_{\mathcal{U}_{0},\mathcal{U}_{L}}-\mathbf{P}_{\mathcal{U}_{0},\mathcal{U}_{L}}\|_{F}/\|\mathbf{P}_{\mathcal{U}_{0},\mathcal{U}_{L}}\|_{F}$. The error rates (averaged over $100$ independent Monte Carlo replicates) for different algorithms are presented in Figure 5, and it shows that CMMI outperforms all competing methods in terms of recovery accuracy. CMMI is also computationally efficient; see Section B.2 of the supplementary material for details.